OpenStudy (anonymous):

Suppose I have a population with μ = 25 and σ2=16. I take 10 samples of size 40. a. What would be your best guess for the value of each of the sample means? (0.25 point)

OpenStudy (anonymous):

Please show work as well.

OpenStudy (amistre64):

im sure theres a formula that can go along with this ....

OpenStudy (amistre64):

http://www.stat.yale.edu/Courses/1997-98/101/sampmn.htm looks like theres someting in here that might be useful

OpenStudy (anonymous):

I've already been to that website but I can't make sense of it.

OpenStudy (amistre64):

is the a sigma^2? or a sigma_2?

OpenStudy (anonymous):

it's to the second power

OpenStudy (amistre64):

off hand id say my best guess of each sample mean is about 25 but i cant confirm that yet, still trying to read up on it

OpenStudy (amistre64):

the formula with all our parts is: \[\frac{\bar x -\mu}{\sigma/\sqrt{n}}\] but im not sure what the "equals" to yes

OpenStudy (amistre64):

*yet

OpenStudy (anonymous):

well i just need to figure out how to calculate the independent random variables but I dunno how to do that.

OpenStudy (amistre64):

the mean of the means is are the independant random variables in this case

OpenStudy (amistre64):

zarkon is here tho so this might be easy for him :)

OpenStudy (zarkon):

as amistre said....the best guess for each of the means (a point estimate) is 25

OpenStudy (zarkon):

no formula needed

OpenStudy (amistre64):

woohoo!!

OpenStudy (anonymous):

I'm curious as to how to calculate the independent random variables exactly

OpenStudy (amistre64):

using the central limit thrms formula that is an adaptation of the regular z= stuff id assume

OpenStudy (anonymous):

?

OpenStudy (amistre64):

recall:\[z=\frac{x-\mu}{\sigma} \]

OpenStudy (amistre64):

in this case we have to "adjust" the sigma for some reason i cant remember to get:\[z=\frac{\bar x -\mu}{\sigma/\sqrt{n}}\]

OpenStudy (amistre64):

spose you take a sample that has a mean of 33; how close would this be to the actual mean? 33-25 -------- = test statistic (zscore) 4/sqrt{40}